OpenStudy (anonymous):

Find the slope of the line containing the points -5, 1/2 and -3, 1/2

OpenStudy (a_clan):

given (x1,y1) and (x2,y2) y2-y1 slope = ----- x2-x1

OpenStudy (callisto):

slope = (1/2-1/2) / [-5-(-3)] = 0